OpenStudy (anonymous):

Thirty is 12% of what number? A. 360 B. 250 C. 3.6 D. 2.5

OpenStudy (anonymous):

You take the 30 and multiply it by 12% as a decimal to get your answer

OpenStudy (anonymous):

250

OpenStudy (anonymous):

30 * 100/12 = 250

OpenStudy (anonymous):

360?

OpenStudy (anonymous):

360 * 0.12 is 43.2

OpenStudy (anonymous):

Yes, @katrinamarie0803 is right

OpenStudy (anonymous):

The answer is 30 / 0.12

OpenStudy (anonymous):

To do this kind of problem, I would set it up as a proportion and solve

OpenStudy (anonymous):

where you have 30/x = 12/100

OpenStudy (anonymous):

Since 250 * 0.12 would get me 12% of that number, which is 30, I did the opposite and divided 0.12 to get the original number.

OpenStudy (anonymous):

Divided by 0.12*

OpenStudy (anonymous):

The way that you should actually solve it is to have a proportion so you get 12x = 3000 and then divide both sides by 12 to get your number

OpenStudy (anonymous):

I am sure there are other ways to solve it but that is how I remember it in school